The joint probability density function of two continuous random variables X and Y is


fxy(x,y) = { C, 0 <y < 2.y < x < 4-y 10, otherwise
Find the value of c and the correlation of X and Y.
Consider the same two random variables X and Y in problem [1] with the same joint probability density function. Find the mean value of Y when X<1.
Answer #1

Yondom lovciable X andy Sol! Here we have given point of g turo Continuous ;O{y <2, J&X {9-4 lo, ow Now we chow cria y owe pcou (x,y) = E(X4) - E(X) E(0)| So JE (XY) = s ry fry 3,y) dxdy - o 2 9-1 ५ (4-29ीव५ way of drdy = $*** [ca-y)-yol day [(4-y+Yfcm-For dy = Da con = 22 O<<<2 11 0 Similarly fy (%) = -4-8 & Fru (nychcz I (4-24) కార = 1- ghi 0422 jow Now E(X)= 5 * fecondこ こ Corr. (X14) = 2/3 2 x 43 x 3 J13 J8/g 2 x 252 = 11.5 onl corr(x,y) = 1 I ole; there is a perfect Corvelation between x an3 Т- 2 ## - [t] = RPG R - 4), - fpts-) Л 3 2 2 Rр xptx = (1>Х5) 3 E